What is the equation for work done in a reversible isothermal process (processes 1 to 3 in this case)? My course tutor wrote down (P3V3)ln(V1/V3), what is this?
1
Expert's answer
2011-06-15T07:21:03-0400
The expression of work for that process
A=-nRT*ln(V3/V1)
where V1-the initial volume, V3 -the final volume.
From the equation of state, if our substance is a gas we can get an approximate value of the expansion work using the ideal gas equation of state
pV=nRT
Since T=const, we can say that nRT=p3V3
also from the properties of ln-function
ln(V3/V1)=-ln(V1/V3)
So we have
A=p3V3*ln(V1/V3) that's exactly what ur teacher wrote.